How man grams of magnesium cyanide are needed to make 275 mL of a 0.075 M solution?
1
Expert's answer
2020-12-07T03:21:19-0500

n(Mg(CN)2) = V*c = (0.275 L)*(0.075 M) = 0.020625 mol

m(Mg(CN)2) = Mr(Mg(CN)2)*n(Mg(CN)2) = (76.3398 g/mol)*(0.020625 mol) = 1.5745 g.